# Cash Flow Underwriting with Bank Transaction Data: Advancing MSME Financial Inclusion in Malaysia

### v2

@@
Abstract.
Despite accounting for 96.1% of all businesses in Malaysia (Department of Statistics Malaysia, 2025), access to financing remains one of the most persistent challenges faced by Micro, Small, and Medium Enterprises (MSMEs). Newly established or young businesses are often excluded from formal credit markets as traditional underwriting approaches rely heavily on credit bureau data. This study investigates the potential of bank statement data as an alternative data source for credit assessment to promote financial inclusion in emerging markets. Firstly, we propose a cash flow-based underwriting pipeline where we utilise bank statement data for end-to-end data extraction and machine learning credit scoring. Secondly, we introduce a novel dataset of 611 loan applicants from a Malaysian lending institution. Thirdly, we develop and evaluate credit scoring models based on application information and bank transaction-derived features. Empirical results show that the use of such data boosts the performance of all models on our dataset, which can improve credit scoring for new-to-lending MSMEs. Lastly, we intend to release the anonymised bank transaction dataset to facilitate further research on MSMEs financial inclusion within Malaysia’s emerging economy.
Bank Statement Data, Alternative Data, Credit Scoring
copyright: rightsretainedconference: AI for Financial Inclusion, Risk Modeling and Resilience in Emerging Markets Workshop at ACM ICAIF 2025; 16/11/2025; Singaporeccs: Computing methodologies Classification and regression trees
Figure 1. Proposed end-to-end cash flow underwriting workflow leveraging bank statement transaction data for credit scoring.:
1. Introduction
Financial inclusion remains a pressing challenge in emerging markets, where a significant portion of the population lacks sufficient credit history to access formal lending. According to the Securities Commission Malaysia (Malaysia, 2025), MSMEs are estimated to contribute around 60% of Malaysia’s gross domestic product. Despite their significant contributions, MSMEs remain underserved by financial institutions in terms of access to financing, giving rise to an estimated MYR 90 billion funding gap (Corporation, 2017).
One fundamental issue is that many MSMEs do not have lending history. Traditional credit assessment relies heavily on credit bureau data such as repayment history, outstanding obligations, and past delinquencies. While this model works well for established businesses, it has shortcomings for MSMEs with thin credit files:
@@ 2. $i$
    It overlooks real-time financial signals or operational dynamics that more accurately reflect a firm’s present financial health.
@@
    It omits alternative indicators of creditworthiness, such as cash flow consistency, receivables and payables patterns, digital transactions, and behavioural metrics from financial activity.
@@
  • To introduce the first ever Malaysian bank statement dataset based on loan application of MSMEs.
@@ 2. Related Works
With the recent technological advancements in machine learning algorithms, the research community begins to realize the limitations of existing traditional credit scoring methods (Gote and Mendhe, 2024). Such studies are important to enhance financial inclusion of MSMEs in emerging markets where MSMEs are often the backbone of economic growth although they have limited access to financing (Elebe and Imediegwu, 2021). As such, alternative data from non-traditional sources that are not included in standard credit bureau files are being used for credit scoring (Group, 2024).
2.1. Limitations of Traditional Credit Scoring
@@
Traditional credit scoring models used by financial institutions primarily rely on credit bureau data such as payment history, amount of debt, and other indicators (Shivhare, 2024). Unfortunately, reliance on such data creates a high entry barrier for MSMEs in emerging economies where they usually lack audited financial statements or loan servicing history required by financial institutions (Courchane and Baines, 2020; Elebe and Imediegwu, 2021). This leads MSMEs to be perceived as high-risk applicants and results in a perpetual cycle of financial exclusion that severely limits their growth potential (Courchane and Baines, 2020). Besides, traditional methods fail to capture recent cash flow status of MSMEs, which can be an accurate representation of financial health (Elebe and Imediegwu, 2021). A recent survey indicates that 58% of lending institutions feel less confident to make decisions based on traditional credit data only (Watterson, 2024). The increasing dissatisfaction motivates the urgent need to use alternative data for a more adaptive and inclusive credit assessment (Gote and Mendhe, 2024).
2.2. Credit Scoring with Alternative Data
@@
In order to overcome the limitations of traditional methods, credit underwriting using alternative data has emerged as a new approach (Gote and Mendhe, 2024; Watterson, 2024; Ngwenya, 2024). For instance, mobile network data are used for credit scoring in Africa (Ngwenya, 2024; Gathu, 2020) and bank account transactions are used for cash flow underwriting in consumer lending (Watterson, 2024). Recent studies suggest that incorporating transactional data can improve predictive accuracy and expand credit to underserved populations (Djeundje et al., 2021). For instance, research has shown that retail transaction data can help construct alternative credit scores (Lee et al., ). A study examining two Indian financial technology (FinTech) firms found that transactional data from platform activity worked as well as credit bureau data in predicting creditworthiness and improved the predictive accuracy when combined with bureau data (Caire and Vidal, 2024). However, the adoption of alternative data in MSMEs lending in Malaysia remains at an early stage, with limited research focus from the industry.
2.3. Machine Learning Models in Credit Scoring
@@
Machine learning (ML) models have proven to be highly effective in processing financial data to produce accurate credit scores (Bücker et al., 2022; Gunnarsson et al., 2021; Trivedi, 2020). These include statistical models like Logistic Regression, Naive Bayes (NB), and ensemble methods like Random Forest, AdaBoost, and Gradient Boosting (Bücker et al., 2022; Trivedi, 2020). For example, a Random Forest model achieved the best performance on Taiwanese credit card data (Abbas and Hussein, 2024). A NB classifier performed well on a highly imbalanced dataset from a New Zealand lender, especially when undeclared features derived from bank statements were incorporated to supplement application form data (Bunker et al., 2016). Given the positive impact of ML models in credit scoring (Bello, 2024), we aim to improve financial inclusion of Malaysian MSMEs through our ML-based cash flow underwriting workflow.
3. Bank Statement Cash Flow Underwriting
@@
The proposed end-to-end cash flow underwriting workflow enables financial institutions to integrate bank statement transaction data analysis into their credit decision-making process. Its primary objective is to enhance traditional credit underwriting by leveraging bank statement-derived features. As illustrated in Figure 1, the workflow consists of three main layers:
@@
(i) Web Layer (Customer Onboarding): This layer acts as the entry point for all loan application-related data including bank statements. A web-based platform is built to allow MSME owners to submit loan applications and for bank officers to review them.
@@
(ii) Application Layer (Bank Statement Analyser): This layer processes uploaded bank statements from the web interface. Multiple specialized engines automate the extraction and conversion of unstructured transactions into structured data, analyses cash flow-based indicators and bank statement related features, and applies rule-based checks to detect potential fraud and data abnormalities.
@@
(iii) Data & Scoring Layer (Cash Flow Underwriting): This layer stores the analysed data and engineered features in a secure database to ensure traceability, regulatory compliance, and auditability. The stored data are then preprocessed, and feature selection methods are applied to retain the most informative variables. Using the selected features, a predictive model is trained to estimate the probability of default and classify credit risk.
@@
4. Bank Statement Transaction Dataset
Table 1. Statistics of MSME loan application datasets.
@@ 4. Bank Statement Transaction Dataset
This study adopts the Cross-Industry Standard Process for Data Mining (CRISP–DM) framework (Chapman et al., 2000) as the methodological basis. CRISP–DM is a widely recognized process model for systematic FinTech studies (Cheng, 2023; Rawat, 2023; da Rocha and de Sousa Junior, 2010) and comprises six main phases i.e. business understanding, data understanding, data preparation, modeling, evaluation, and deployment. The first phase involves evaluating bank-statement transaction data as an alternative source for MSME credit risk assessment in Malaysia. The second phase focuses on constructing the proposed dataset of 611 MSME loan applicants and is split into training and validation sets with a 60:40 ratio as shown in Table 1. Among these, 518 applicants have a good credit record, while 93 have a history of default. Each applicant’s record contains two main components: application form information submitted during the loan application process (e.g., demographic and business characteristics) and bank statement transaction data capturing detailed inflows and outflows over a six-month period. The third phase is about data preparation and includes several key tasks:
@@
    Data cleaning and de-duplication to ensure consistency.
@@ 2. $i$
    Fixing missing values, standardizing transaction categories and validating data integrity across all records.
@@
    Deriving features and variables from transaction data, such as determining cash flow stability, deposit regularity, and balance volatility. All personally identifiable information was anonymised prior to analysis. To preserve confidentiality, the features are grouped into application form information (7 features) and bank statement features (10 features). Although the specific feature calculations cannot be disclosed due to a non-disclosure agreement, all bank statement features are derived solely from bank statement data.
5. Transaction Data-based Credit Scoring
The fourth phase of CRISP–DM in this study employs Logistic Regression (LR) (Cox, 1958) on application form and bank transaction data derived from our dataset as the baseline credit scoring model. LR is widely adopted in credit risk modeling due to its interpretability, statistical robustness, and capacity to estimate the probability of default (Bücker et al., 2022; Trivedi, 2020; Abbas and Hussein, 2024; Bunker et al., 2016). To ensure that only informative predictors are retained, feature selection and transformation are guided by the Weight of Evidence (WOE) and Information Value (IV) framework, which are widely used in credit risk modeling for quantifying the predictive power of individual variables (Ngwenya, 2024). In practice, each feature is divided into discrete intervals based on thresholds derived from the data. These bins allow the calculation of WOE and IV at the group level, which makes them easier to interpret within a logistic regression model. The following notation is used to quantify the distribution of defaults and non-defaults across feature bins:
@@ 5. Transaction Data-based Credit Scoring
(3) $\text{Dist}^{(b)}_{jk}$ $=\frac{n_{bjk}}{N_{b}},$ $\text{Dist}^{(g)}_{jk}$ $=\frac{n_{gjk}}{N_{g}}.$
5.1. Weight of Evidence (WOE)
@@
(4) $\mathrm{WOE}_{jk}=\log\!\left(\frac{\text{Dist}^{(g)}_{jk}}{\text{Dist}^{(b)}_{jk}}\right)=\log\!\left(\frac{n_{gjk}/N_{g}}{n_{bjk}/N_{b}}\right).$
A positive $\mathrm{WOE}_{jk}$ indicates that bin $B_{jk}$ is more common among non-default cases, suggesting lower risk, whereas a negative value indicates a higher likelihood of default. Besides, a small smoothing constant $\epsilon>0$ is added to both numerators and denominators to handle bins with zero counts. Once WOE is computed for each bin, the predictive strength of feature $j$ can be summarized by its IV.
5.2. Information Value (IV)
@@
(5) $\mathrm{IV}_{j}=\sum_{k=1}^{K_{j}}\left(\text{Dist}^{(g)}_{jk}-\text{Dist}^{(b)}_{jk}\right)\mathrm{WOE}_{jk}.$
A larger $\mathrm{IV}_{j}$ indicates stronger discriminatory power of feature $j$ between default and non-default classes. In this study, IV serves as both a feature-ranking criterion and an interpretability measure, helping to identify the most influential transaction-level indicators for creditworthiness. Following industry practice (Siddiqi, 2017), we adopt the commonly used thresholds: $\mathrm{IV}<0.02$ (not predictive), $0.02\leq\mathrm{IV}<0.1$ (weak), $0.1\leq\mathrm{IV}<0.3$ (medium), $0.3\leq\mathrm{IV}<0.5$ (strong), and $\mathrm{IV}\geq 0.5$ (suspiciously high; suggesting potential data leakage).
5.3. Binning of Bank Statement Features
Specifically, we use quantile or supervised monotonic binning to obtain $\{B_{jk}\}_{k=1}^{K_{j}}$ for continuous transaction-derived variables (e.g., cash-flow stability, balance volatility). Then, we group rare datapoints to ensure each bin has sufficient defaults and non-defaults for sparse categorical variables (e.g., business sector code). Computations of WOE and IV are performed on training folds only to avoid data leakage.
5.4. Logistic Regression Model
@@
(6) $\log\frac{P(y_{i}=1\,|\,\mathbf{x}_{i})}{P(y_{i}=0\,|\,\mathbf{x}_{i})}=\beta_{0}+\sum_{j=1}^{d}\beta_{j}\,\mathrm{WOE}_{j}(x_{ij}).$
where $\beta_{0}$ is the intercept and $\beta_{j}$ represents the coefficient for feature $j$. This allows direct interpretation of coefficients in terms of the credit risk associated with each feature. Positive $\beta_{j}$ values indicate that higher WOE corresponds to lower default risk, and vice versa. Hence, WOE encoding stabilizes estimation and supports consistent, monotonic relationships between predictors and credit outcomes.
@@
(7) $P(y_{i}=1\mid\mathbf{x}_{i};\boldsymbol{\beta})=\sigma(\beta_{0}+\mathbf{x}_{i}^{\top}\boldsymbol{\beta}),$
@@
(8) $\mathcal{L}(\boldsymbol{\beta})=\sum_{i=1}^{n}\left[y_{i}\log p_{i}+(1-y_{i})\log(1-p_{i})\right]-\lambda\,\|\boldsymbol{\beta}\|_{2}^{2},$
where $p_{i}=P(y_{i}=1\mid\mathbf{x}_{i};\boldsymbol{\beta})$ and $\lambda\geq 0$ controls the strength of the $\ell_{2}$ regularization (Ng, 2004). This mitigates overfitting by shrinking coefficient magnitudes, which is particularly important when modeling high-dimensional, transaction-derived features.
Figure 2. ROC curves of all models evaluated on the validation split of the proposed dataset.:
Figure 3. Information Value (IV) distribution of features derived from application and bank statement data. Features with higher IV indicate stronger predictive power for creditworthiness classification.:
6. Experiments and Deployment
Aligned with the fifth phase of CRISP–DM, we conduct a series of experiments to evaluate the effectiveness of the proposed cash flow underwriting workflow and the role of bank statement data in enhancing MSME credit scoring performance. The experiments include comparative analyses with established machine learning methods and ablation studies on transaction-derived features.
6.1. Implementation Details
We benchmark the baseline LR model against several widely used machine learning methods, including Random Forest (RF) (Breiman, 2001), Gradient Boosting (GB) (Friedman, 2001), and AdaBoost (AB) (Schapire, 2013). All models are implemented in scikit-learn (scikit learn, 2025c, a, b, d). For the results in Section 6.2, models were trained with default hyperparameters. In Section 6.3, hyperparameters were tuned using a randomized grid search with 50 trials to ensure robustness. Model performance is evaluated using the Area Under the Receiver Operating Characteristic Curve (AUROC) (Bradley, 1997), which measures the ability to discriminate between default and non-default cases across varying thresholds. An AUROC of 0.5 indicates no discriminative power (equivalent to random guessing), whereas a value of 1.0 indicates perfect discrimination. To ensure interpretability and systematic analysis, both application and bank statement features are grouped into three categories:

### v3

@@
Abstract
Despite accounting for 96.1% of all businesses in Malaysia (smeprofile), access to financing remains one of the most persistent challenges faced by Micro, Small, and Medium Enterprises (MSMEs). Newly established businesses are often excluded from formal credit markets as traditional underwriting approaches rely heavily on credit bureau data. This study investigates the potential of bank statement data as an alternative data source for credit assessment to promote financial inclusion in emerging markets. First, we propose a cash flow-based underwriting pipeline where we utilise bank statement data for end-to-end data extraction and machine learning credit scoring. Second, we introduce a novel dataset of 611 loan applicants from a Malaysian lending institution. Third, we develop and evaluate credit scoring models based on application information and bank transaction-derived features. Empirical results show that the use of such data boosts the performance of all models on our dataset, which can improve credit scoring for new-to-lending MSMEs. Finally, we will release the anonymised bank transaction dataset to facilitate further research on MSME financial inclusion within Malaysia’s emerging economy.
1 Introduction
Figure 1: Proposed end-to-end cash flow underwriting workflow leveraging bank statement transaction data for credit scoring.
Financial inclusion remains a pressing challenge in emerging markets, where a significant portion of the population lacks sufficient credit history to access formal lending. According to the Securities Commission Malaysia (cm_plan), MSMEs are estimated to contribute around 60% of Malaysia’s gross domestic product. Despite their significant contributions, MSMEs remain underserved by financial institutions in terms of access to financing, giving rise to an estimated MYR 90 billion funding gap (international2017msme).
One fundamental issue is that many MSMEs do not have lending history. Traditional credit assessment relies heavily on credit bureau data such as repayment history, outstanding obligations, and past delinquencies. While this traditional credit model works well for established businesses, it has shortcomings for MSMEs with thin credit files:
@@ 2. $i$
    It overlooks real-time financial signals or operational dynamics that could more accurately reflect a firm’s present financial health.
@@
    It omits alternative indicators of creditworthiness, such as cash flow consistency, receivables and payables patterns, digital transactions, and behavioural metrics from daily financial activity.
@@
  • To introduce the first ever Malaysian bank statement dataset based on loan applications of MSMEs.
@@ 2 Related Work
With the recent technological advancements in machine learning algorithms, the research community has recognised the limitations of existing traditional credit scoring methods (gote2024building). Such studies are crucial for enhancing financial inclusion of MSMEs in emerging markets, where MSMEs are often the backbone of economic growth despite having limited access to financing (elebe2021credit). As such, alternative data from non-traditional sources that are not included in standard credit bureau files are being used for credit scoring (world_bank).
Limitations of Traditional Credit Scoring
@@
Traditional credit scoring models used by financial institutions primarily rely on credit bureau data such as payment history, amount of debt, and other indicators (shivhare2024utilize). Unfortunately, reliance on such data creates a high entry barrier for MSMEs in emerging economies where they usually lack audited financial statements or loan servicing history required by financial institutions (courchane2019use; elebe2021credit). This leads MSMEs to be perceived as high-risk applicants and results in a perpetual cycle of financial exclusion that severely limits their growth potential (courchane2019use). Besides, traditional methods fail to capture recent cash flow status of MSMEs, which can be an accurate representation of financial health (elebe2021credit). A recent survey indicates that 58% of lending institutions feel less confident to make decisions based on traditional credit data only (datos). The increasing dissatisfaction motivates the urgent need to use alternative data for a more adaptive and inclusive credit assessment (gote2024building).
Credit Scoring with Alternative Data
@@
In order to overcome the limitations of traditional methods, credit underwriting using alternative data has emerged as a new approach (gote2024building; datos; africa). For instance, mobile network data are used for credit scoring in Africa (africa; kenya) and bank account transactions are used for cash flow underwriting in consumer lending (datos). Recent studies suggest that incorporating transactional data can improve predictive accuracy and expand credit to underserved populations (djeundje2021enhancing). For instance, research has shown that retail transaction data can help construct alternative credit scores (express_peru). A study examining two Indian financial technology (FinTech) firms found that transactional data from platform activity worked as well as credit bureau data in predicting creditworthiness and improved the predictive accuracy when combined with bureau data (cgap). However, the adoption of alternative data in MSME lending in Malaysia remains at an early stage, with limited research focus from the industry.
Machine Learning Models in Credit Scoring
@@
Machine learning (ML) models have proven to be highly effective in processing financial data to produce accurate credit scores (bucker2022transparency; gunnarsson2021deep; trivedi2020study). These include statistical models like Logistic Regression, Naive Bayes (NB), and ensemble methods like Random Forest, AdaBoost, and Gradient Boosting (bucker2022transparency; trivedi2020study). For example, a Random Forest model achieved the best performance on Taiwanese credit card data (abbas2024algorithm). A NB classifier performed well on a highly imbalanced dataset from a New Zealand lender, especially when undeclared features derived from bank statements were incorporated to supplement application form data (bunker2016improving). Given the positive impact of ML models in credit scoring, we aim to improve financial inclusion of Malaysian MSMEs through our ML-based cash flow underwriting workflow.
Figure 2: Overview of AI modules with their objectives and example outputs within the Bank Statement Analyser.
Agentic Workflow in Credit Scoring
The deployment of Large Language Model (LLM)-powered agentic systems marks a major advancement in automated credit scoring, loan approval, and customer risk profiling within financial services (okpala2025agentic; ali2025efficient; paleti2024agentic). Recent agentic frameworks have demonstrated strong performance in complex model risk management tasks, including credit card approval and portfolio risk modelling (okpala2025agentic). Beyond these use cases, agentic workflows are transforming banking by enabling adaptive credit profiling and predictive loan approvals, thereby improving risk assessment accuracy through bias detection and intelligent decision automation (paleti2024agentic). Similar approaches have been explored in mortgage lending, where machine learning–driven agents streamline underwriting and accelerate credit evaluation (chitturi2025agentic). However, existing methods remain limited in production deployment and lack a fully grounded, end-to-end workflow for MSMEs credit scoring.
3 Bank Statement Cash Flow Underwriting
@@
The proposed end-to-end cash flow underwriting workflow enables financial institutions to integrate bank statement transaction data analysis into their credit decision-making process. Its primary objective is to enhance traditional credit underwriting by leveraging bank statement-derived features. As illustrated in Figure 1, the workflow consists of three main layers, which serve different purposes.
Web Layer (Customer Onboarding)
@@
This layer acts as the entry point for all loan application-related data including bank statements. A web-based platform is built to allow MSME owners to submit loan applications and for bank officers to review them.
Application Layer (Bank Statement Analyser)
@@
This layer processes uploaded bank statements from the web interface. Multiple specialised engines automate the extraction and conversion of unstructured transactions into structured data, analyse cash flow-based indicators and bank statement related features, and apply rule-based checks to detect potential fraud and data abnormalities. Within these extraction and analytic engines, several Artificial Intelligence (AI) modules perform complex and dedicated tasks to derive actionable insights from the uploaded bank statements as depicted in Figure 2. The outputs of these modules form the foundation for downstream credit scoring.
Key Information Extraction Module
The key information extraction module employs Optical Character Recognition (OCR) to extract essential fields from bank statements, including the account number, account holder name, address, and statement date. These attributes are cross validated across statements to ensure accurate document ownership verification. These fields enable the engine to link multiple accounts belonging to the same business entity, providing a holistic view of its financial activity across accounts.
Transaction Table Extraction Module
The transaction table extraction module localizes and digitizes tabular transaction data using OCR and layout analysis. It identifies table headers, then maps rows and columns to extract transaction date, description, debit, credit, and balance amounts. Additional pre-processing handles merged cells, page breaks, and format inconsistencies to preserve data fidelity. The resulting structured transaction records serve as critical inputs for downstream analytical modules that evaluate transaction patterns and account-level cash flow dynamics.
Fraud Analysis Module
The fraud analysis module leverages Computer Vision and rule-based techniques to detect tampered or falsified bank statements. For each document, it examines both visual and metadata cues such as inconsistent font styles or sizes, altered layouts, metadata mismatches, and abnormal pixel-level patterns that may indicate digital editing. Once potential tampering is identified, the affected statements are automatically flagged and returned to bank officers for manual verification. This module provides an additional layer of security and assurance within the underwriting workflow, ensuring that only authentic bank statements are analysed for credit scoring.
Network Analysis Module
The network analysis module constructs transaction networks from the extracted transaction data to uncover relationships among senders and beneficiaries across multiple bank accounts. Using graph-based algorithms, it maps transaction linkages to detect circular fund flows, kiting activities, and transfers to blacklisted or interrelated entities. This module also quantifies network connectivity and transaction density to reveal abnormal money movement patterns that may indicate fraudulent activity or financial distress, thereby enhancing overall risk assessment.
Cash Flow Analysis Module
The cash flow analysis module evaluates cash flow status based on the inflows and outflows extracted from transaction records. It then applies Natural Language Processing to infer transaction intent and classify entries into meaningful categories based on their descriptions. Using these classifications, this module groups related transactions to identify spending and receipt patterns, enabling the detection of outliers that deviate from typical cash flow behaviour. It also computes key metrics such as average, highest, and lowest balances, quantifying cash flow health and provides predictive features for credit scoring.
Data & Scoring Layer (Cash Flow Underwriting)
@@
This layer stores the analysed data and engineered features in a secure database to ensure traceability, regulatory compliance, and auditability. The stored data are then preprocessed, and feature selection methods are applied to retain the most informative variables. Using the selected features, a predictive model is trained to estimate the probability of default and classify credit risk.
@@
4 Bank Statement Transaction Dataset
Table 1: Statistics of MSME loan application datasets.
@@ 4 Bank Statement Transaction Dataset
This study adopts the Cross-Industry Standard Process for Data Mining (CRISP–DM) framework (crisp_dm) as the methodological basis. CRISP–DM is a widely recognised process model for systematic FinTech studies (cheng2023evaluating; rawat2023applying; da2010identifying) and comprises six main phases i.e. business understanding, data understanding, data preparation, modelling, evaluation, and deployment. The first phase involves evaluating bank statement transaction data as an alternative source for MSME credit risk assessment in Malaysia. The second phase focuses on constructing the proposed dataset of 611 MSME loan applicants and is split into training and validation sets with a 60:40 ratio as shown in Table 1. Among these, 518 applicants have a good credit record, while 93 have a history of default. Each applicant’s record contains two main components: application form information submitted during the loan application process (e.g., demographic and business characteristics) and bank statement transaction data capturing detailed inflows and outflows over a six-month period. The third phase involves data preparation and includes several key tasks:
@@
    Cleaning and de-duplicating data to ensure consistency.
@@ 2. $i$
    Fixing missing values, standardising transaction categories, and validating data integrity across all records.
@@
    Deriving features and variables from transaction data, such as determining cash flow stability, deposit regularity, and balance volatility.
All personally identifiable information was anonymised prior to analysis. To preserve confidentiality, the features are grouped into application form information (7 features) and bank statement features (10 features). Although the specific feature calculations cannot be disclosed due to a non-disclosure agreement, all bank statement features are derived solely from bank statement data.
5 Transaction Data-based Credit Scoring
The fourth phase of CRISP–DM in this study employs Logistic Regression (LR) (cox1958regression) on application form and bank transaction data derived from our dataset as the baseline credit scoring model. LR is widely adopted in credit risk modelling due to its interpretability, statistical robustness, and capacity to estimate the probability of default (bucker2022transparency; trivedi2020study; abbas2024algorithm; bunker2016improving). To ensure that only informative predictors are retained, feature selection and transformation are guided by the Weight of Evidence (WOE) and Information Value (IV) framework, which are widely used in credit risk modelling for quantifying the predictive power of individual variables (africa). In practice, each feature is divided into discrete intervals based on suitable thresholds determined from the data. These bins allow the calculation of WOE and IV at the group level, which makes them easier to interpret within a logistic regression model. The following notation is used to quantify the distribution of defaults and non-defaults across feature bins:
@@ 5 Transaction Data-based Credit Scoring
$\text{Dist}^{(b)}_{jk}$ $=\frac{n_{bjk}}{N_{b}},$ $\text{Dist}^{(g)}_{jk}$ $=\frac{n_{gjk}}{N_{g}}.$ (3)
Weight of Evidence (WOE)
@@
$\mathrm{WOE}_{jk}=\log\!\left(\frac{\text{Dist}^{(g)}_{jk}}{\text{Dist}^{(b)}_{jk}}\right)=\log\!\left(\frac{n_{gjk}/N_{g}}{n_{bjk}/N_{b}}\right).$ (4)
A positive $\mathrm{WOE}_{jk}$ indicates that bin $B_{jk}$ is more common among non-default cases, suggesting lower risk, whereas a negative value indicates a higher likelihood of default. Once WOE is computed for each bin, the overall predictive strength of feature $j$ can then be summarised by its IV.
Information Value (IV)
@@
$\mathrm{IV}_{j}=\sum_{k=1}^{K_{j}}\left(\text{Dist}^{(g)}_{jk}-\text{Dist}^{(b)}_{jk}\right)\mathrm{WOE}_{jk}.$ (5)
A larger $\mathrm{IV}_{j}$ indicates stronger discriminatory power of feature $j$ between default and non-default classes. In this study, IV serves as both a feature-ranking criterion and an interpretability measure, helping to identify the most influential transaction-level indicators for creditworthiness. Following industry practice (intel_credit_scoring), we adopt the commonly used interpretive thresholds: $\mathrm{IV}<0.02$ (not predictive), $0.02\leq\mathrm{IV}<0.1$ (weak), $0.1\leq\mathrm{IV}<0.3$ (medium), $0.3\leq\mathrm{IV}<0.5$ (strong), and $\mathrm{IV}\geq 0.5$ (suspiciously high; suggesting potential data leakage).
Binning of Bank Statement Features
Specifically, we use quantile or supervised monotonic binning to obtain $\{B_{jk}\}_{k=1}^{K_{j}}$ for continuous transaction-derived variables (e.g., cash flow stability, balance volatility). Then, we group rare datapoints to ensure each bin has sufficient defaults and non-defaults for sparse categorical variables (e.g., business sector code). Computations of WOE and IV are performed on training folds only to avoid data leakage.
Logistic Regression Model
@@
$\log\frac{P(y_{i}=1\,|\,\mathbf{x}_{i})}{P(y_{i}=0\,|\,\mathbf{x}_{i})}=\beta_{0}+\sum_{j=1}^{d}\beta_{j}\,\mathrm{WOE}_{j}(x_{ij}).$ (6)
where $\beta_{0}$ is the intercept and $\beta_{j}$ represents the coefficient for feature $j$. This allows direct interpretation of coefficients in terms of the credit risk associated with each feature. Positive $\beta_{j}$ values indicate that higher WOE corresponds to lower default risk, and vice versa. Hence, WOE encoding stabilises estimation and supports consistent, monotonic relationships between predictors and credit outcomes.
@@
$P(y_{i}=1\mid\mathbf{x}_{i};\boldsymbol{\beta})=\sigma(\beta_{0}+\mathbf{x}_{i}^{\top}\boldsymbol{\beta}),$ (7)
@@
$\mathcal{L}(\boldsymbol{\beta})=\sum_{i=1}^{n}\left[y_{i}\log p_{i}+(1-y_{i})\log(1-p_{i})\right]-\lambda\,\|\boldsymbol{\beta}\|_{2}^{2},$ (8)
where $p_{i}=P(y_{i}=1\mid\mathbf{x}_{i};\boldsymbol{\beta})$ and $\lambda\geq 0$ controls the strength of the $\ell_{2}$ regularization (ng2004feature). This mitigates overfitting by shrinking coefficient magnitudes, which is particularly important when modelling high-dimensional, transaction-derived features.
6 Experiments and Deployment
Following the fifth phase of CRISP–DM, we conduct a series of experiments to evaluate the effectiveness of the proposed cash flow underwriting workflow and the role of bank statement data in enhancing MSME credit scoring performance. The experiments include comparative analyses with established machine learning methods and detailed ablation studies on transaction-derived features.
Figure 3: ROC curves of all models evaluated on the validation split of the proposed dataset.
(a) IV of features derived from application and bank statement data.
(b) AUROC on train and validation splits with different combinations.
Figure 4: Results of ablation studies on IV and AUROC performance on the proposed dataset. Best viewed in colour.
Implementation Details
We benchmark the baseline LR model against several widely used machine learning methods, including Random Forest (RF) (breiman2001random), Gradient Boosting (GB) (friedman2001greedy), and AdaBoost (AB) (schapire2013explaining). All models are implemented in scikit-learn (sklearn_lr; sklearn_ada; sklearn_gb; sklearn_rf). For the results in Section 6, models were trained with default hyperparameters. In Section 6, hyperparameters were tuned using a randomized grid search with 50 trials to ensure robustness. We evaluate model performance using the Area Under the Receiver Operating Characteristic Curve (AUROC) (bradley1997use), which measures the ability to discriminate between default and non-default cases across varying thresholds. An AUROC of 0.5 indicates no discriminative power (equivalent to random guessing), whereas a value of 1.0 indicates perfect discrimination. To ensure interpretability and systematic analysis, both application and bank statement features are grouped into three categories: